### v1

# Tight bound for the Erdős-Pósa property of tree minors

Abstract.
Let $T$ be a tree on $t$ vertices. We prove that for every positive integer $k$ and every graph $G$, either $G$ contains $k$ pairwise vertex-disjoint subgraphs each having a $T$ minor, or there exists a set $X$ of at most $t(k-1)$ vertices of $G$ such that $G-X$ has no $T$ minor. The bound on the size of $X$ is best possible and improves on an earlier $f(t)k$ bound proved by Fiorini, Joret, and Wood (2013) with some very fast growing function $f(t)$. Moreover, our proof is very short and simple.
G. Joret is supported by a PDR grant from the Belgian National Fund for Scientific Research (FNRS). V. Dujmović is supported by NSERC and a University of Ottawa Research Chair. P. Micek is supported by the National Science Center of Poland under grant UMO-2023/05/Y/ST6/00079 within the WEAVE-UNISONO program. P. Morin is supported by NSERC.
1. Introduction
In 1965, Erdős and Pósa [EP65] showed that every graph $G$ either contains $k$ vertex-disjoint cycles or contains a set $X$ of $\mathcal{O}(k\log k)$ vertices such that $G-X$ has no cycles. The $\mathcal{O}(k\log k)$ bound on the size of $X$ is best possible up to a constant factor. Using their Grid Minor Theorem, Robertson and Seymour [RS86] proved the following generalization: For every planar graph $H$, there exists a function $f_{H}(k)$ such that every graph $G$ contains either $k$ vertex-disjoint subgraphs each having an $H$ minor, or a set $X$ of at most $f_{H}(k)$ vertices such that $G-X$ has no $H$ minor. For $H=K_{3}$, this corresponds to the setting of the Erdős-Pósa theorem.
The theorem of Robertson and Seymour is best possible in the sense that no such result holds when $H$ is not planar. The original upper bound of $f_{H}(k)$ on the size of $X$ depends on bounds from the Grid Minor Theorem and is large as a result (though it is polynomial in $k$ if we use the polynomial version of the Grid Minor Theorem, see [CC16]). Chekuri and Chuzhoy [CC13] subsequently showed an improved upper bound of $\mathcal{O}_{H}(k\log^{c}k)$ for fixed $H$, where $c$ is some large but absolute constant. This was in turn improved to $\mathcal{O}_{H}(k\log k)$ by Cames van Batenburg, Huynh, Joret, and Raymond [CvBHJR19], thus matching the original bound of Erdős and Pósa for cycles.
An $\mathcal{O}_{H}(k\log k)$ bound is best possible when $H$ contains a cycle. However, when $H$ is a forest, it turns out that one can obtain a linear in $k$ bound on the size of $X$, as proved by Fiorini, Joret, and Wood [FJW13]. The original proof is technical and gives an $\mathcal{O}_{H}(k)$ bound with a non-explicit constant factor that grows extremely fast as a function of $|V(H)|$. (This is due to the use of MSO-based tools in the proof, among others.) In this short note we give a very simple proof of their result with an optimal dependence on $t$ and $k$ when $H$ is a tree.
Theorem 1.
Let $T$ be a tree on $t$ vertices. For every positive integer $k$ and every graph $G$, either $G$ contains $k$ pairwise vertex-disjoint subgraphs each having a $T$ minor, or there exists a set $X$ of at most $t(k-1)$ vertices of $G$ such that $G-X$ has no $T$ minor.
Observe that the bound on the size of $X$ in Theorem 1 is tight: If $G$ is a complete graph on $tk-1$ vertices, then $G$ does not contain $k$ pairwise vertex-disjoint subgraphs each having a $T$ minor, and every set $X$ of vertices such that $G-X$ has no $T$ minor has size at least $|V(G)|-(t-1)=t(k-1)$.
Theorem 1 follows immediately from the following more general result for forests.
Theorem 2.
Let $F$ be a forest on $t$ vertices and let $t^{\prime}$ be the maximum number of vertices in a component of $F$. For every positive integer $k$ and every graph $G$, either $G$ contains $k$ pairwise vertex-disjoint subgraphs each having an $F$ minor, or there exists a set $X$ of at most $tk-t^{\prime}$ vertices of $G$ such that $G-X$ has no $F$ minor.
Let us also point out the following corollary of Theorem 1 (proved in the next section).
Corollary 3.
For all positive integers $p$ and $k$, and for every graph $G$, either $G$ contains $k$ vertex-disjoint subgraphs each of pathwidth at least $p$, or $G$ contains a set $X$ of at most $2\cdot 3^{p+1}k$ vertices such that $G-X$ has pathwidth strictly less than $p$.
2. Proof
For a positive integer $k$, we use the notation $[k]\coloneqq\{1,\ldots,k\}$, and when $k=0$ let $[k]\coloneqq\varnothing$.
Let $G$ be a graph. We denote by $V(G)$ and $E(G)$, the vertex set and edge set of $G$, respectively. Let $X\subseteq V(G)$. Then $G[X]$ denotes the subgraph of $G$ induced by the vertices in $X$ and $G-X=G[V(G)-X]$. We define the boundary of $X$ in $G$ to be $\partial_{G}X:=\{v\in X\mid vw\in E(G),\,w\in V(G-X)\}$. We omit the subscript $G$ when the graph $G$ is clear from the context.
A path decomposition of $G$ is a sequence $(B_{1},B_{2},\dots,B_{q})$ of vertex subsets of $G$ called bags satisfying the following properties: (1) every vertex of $G$ appears in a nonempty set of consecutive bags, and (2) for every edge $uv$ of $G$, there is a bag containing both $u$ and $v$. The width of the path decomposition is the maximum size of a bag minus $1$. The pathwidth $\operatorname{pw}(G)$ of $G$ is the minimum width of a path decomposition of $G$.
A graph $H$ is a minor of a graph $G$ if $H$ can be obtained from a subgraph of $G$ by contracting edges. Robertson and Seymour [RS83] proved that there exists a function $f:\mathbb{N}\to\mathbb{N}$ such that for every graph $G$ and every forest $F$ on $t$ vertices, if $\operatorname{pw}(G)\geqslant f(t)$ then $G$ contains $F$ as a minor. Bienstock, Robertson, Seymour, and Thomas [BRST91] later showed that one can take $f(t)=t-1$, which is best possible. Diestel [Die95] subsequently gave a short proof of this result. Our proof of Theorem 2 builds on the following slightly stronger result, which appears implicitly in Diestel’s proof [Die95].
Lemma 4 ([Die95]).
Let $G$ be a graph, let $t$ be a positive integer, and let $F$ be a forest on $t$ vertices. If $\operatorname{pw}(G)\geqslant t-1$, then there exists $Y\subseteq V(G)$ such that
  1. (1)
    $G[Y]$ has a path decomposition $(B_{1},\ldots,B_{q})$ of width at most $t-1$ such that $\partial Y\subseteq B_{q}$, and
  2. (2)
    $G[Y]$ contains $F$ as a minor.
We now turn to the proof of Theorem 2.
Proof of Theorem 2.
We prove the following strengthening of Theorem 2: Let $G$ be a graph, let $c$ be a positive integer, let $t_{1}\leqslant\cdots\leqslant t_{c}$ be non-negative integers, let $T_{1},\ldots,T_{c}$ be trees with $|V(T_{i})|=t_{i}$ for every $i\in[c]$, let $x_{1},\ldots,x_{c}$ be non-negative integers, at least one of which is nonzero, and let $I:=\{i\in[c]:x_{i}\geqslant 1\}$. Then either
  1. (1)
    $G$ contains pairwise vertex-disjoint subgraphs $\{M_{i,j}:i\in[c],\,j\in[x_{i}]\}$ such that, for each $i\in[c]$ and $j\in[x_{i}]$, $M_{i,j}$ contains a $T_{i}$ minor, or
  2. (2)
    there exists $X\subseteq V(G)$ with $|X|\leqslant\sum_{i\in I}x_{i}t_{i}-t_{\max(I)}$ and $G-X$ does not contain $T_{i}$ as a minor for some $i\in I$.
We call the tuple $(G,c,T_{1},\ldots,T_{c},x_{1},\ldots,x_{c})$ an instance. Theorem 2 follows by letting $T_{1},\ldots,T_{c}$ be the components of the forest $F$ and letting $x_{1}=x_{2}=\cdots=x_{c}=k$.
Let $(G,c,T_{1},\ldots,T_{c},x_{1}\ldots,x_{c})$ be an instance, and let $m=\min(I)$. Then $T_{m}$ is a smallest tree among $T_{1},\ldots,T_{c}$ such that $x_{m}\geqslant 1$, that is, such that we are still missing a model of $T_{m}$. In the base case, $\sum_{i\in[c]}x_{i}=x_{m}=1$, and either $G$ has a $T_{m}$ minor and the first outcome of the statement holds, or $G$ has no such minor and the second outcome holds with $X:=\varnothing$.
For the inductive case, assume that $\sum_{i\in[c]}x_{i}\geqslant 2$, and that the statement holds for instances with smaller values of the sum. If $G$ has pathwidth at least $t_{m}-1$, apply Lemma 4 with $t=t_{m}$ and $F=T_{m}$, and let $Y$ be the resulting subset of vertices of $G$. If $G$ has pathwidth less than $t_{m}-1$, simply let $Y:=V(G)$. In either case, $G[Y]$ has pathwidth at most $t_{m}-1$ and has a path decomposition $(B_{1},B_{2},\dots,B_{q})$ with $|B_{\ell}|\leqslant t_{m}$ for all $\ell\in[q]$, and such that $\partial_{G}Y\subseteq B_{q}$. See Figure 1.
Figure 1. The set $Y$ and the graph $G_{\ell}$ whose boundary in $G$ is contained in $B_{\ell}$.
If, for every $i\in I$, $G[Y]$ has no $T_{i}$ minor then, by Lemma 4, the pathwidth of $G$ is less than $t_{m}-1$ so $Y=V(G)$ and the second outcome of the statement holds with $X:=\varnothing$. So we may assume that $G[Y]$ has a $T_{i}$ minor for some $i\in I$. Let $\ell\in[q]$ be the smallest index such that $G_{\ell}:=G[B_{1}\cup\cdots\cup B_{\ell}]$ contains a $T_{i}$ minor for some $i\in I$ and let $i^{\prime}$ be any index in $I$ such that $G_{\ell}$ contains a $T_{i^{\prime}}$ minor. Observe that
$G_{\ell}-B_{\ell}$ has no $T_{i}$ minor for any $i\in I$. ($\star$)
Also, we claim that there is no edge in $G$ between vertices of $G_{\ell}-B_{\ell}$ and vertices of $G-V(G_{\ell})$. To see this, suppose for a contradiction that $uv$ is such an edge, with $u\in V(G_{\ell})-B_{\ell}$ and $v\in V(G)-V(G_{\ell})$. If $v\in Y$, then $u$ and $v$ appear together in some bag $B_{j}$ of the path decomposition $(B_{1},B_{2},\dots,B_{q})$ of $G[Y]$, and $j>\ell$ since $v\notin B_{1}\cup\cdots\cup B_{\ell}$. However, since $u\in B_{1}\cup\cdots\cup B_{\ell-1}$, this implies that $u$ belongs also to $B_{\ell}$, a contradiction. If $v\notin Y$, then $u\in\partial Y$, and thus $u\in B_{q}$. Again, we deduce that $u\in B_{\ell}$, a contradiction.
Let $G^{\prime}:=G-V(G_{\ell})$. Let $x^{\prime}_{i}:=x_{i}$ for each $i\in[c]-\{i^{\prime}\}$ and let $x^{\prime}_{i^{\prime}}:=x_{i^{\prime}}-1$. Let $I^{\prime}=\{i\in[c]:x^{\prime}_{i}\geqslant 1\}$. Apply induction to the instance $(G^{\prime},c,T_{1},\ldots,T_{c},x^{\prime}_{1},\ldots,x^{\prime}_{c})$. If it results in a set of vertex-disjoint subgraphs $\{M^{\prime}_{i,j}:i\in[c],\,j\in[x^{\prime}_{i}]\}$, with $M_{i,j}$ containing a $T_{i}$ minor for each $i\in[c]$ and $j\in[x^{\prime}_{i}]$, then we let $M_{i,j}:=M^{\prime}_{i,j}$ for each $i\in[c]$ and $j\in[x^{\prime}_{i}]$, and $M_{i^{\prime},x_{i^{\prime}}}:=G_{\ell}$, which results in the desired collection of vertex-disjoint subgraphs. Otherwise, we obtain a set $X^{\prime}$ of at most $\sum_{i\in I^{\prime}}x^{\prime}_{i}t_{i}-t_{\max(I^{\prime})}$ vertices such that $G^{\prime}-X^{\prime}$ does not contain $T_{a}$ as a minor for some $a\in I^{\prime}$.
Let $X:=X^{\prime}\cup B_{\ell}$. Observe that
$|X|=|X^{\prime}|+|B_{\ell}|$ $\leqslant\sum_{i\in I^{\prime}}x^{\prime}_{i}t_{i}-t_{\max(I^{\prime})}+t_{m}\leqslant\sum_{i\in I}x_{i}t_{i}-(t_{\max(I^{\prime})}+t_{i^{\prime}}-t_{m})$
$\leqslant\sum_{i\in I}x_{i}t_{i}-t_{\max(I)}.$
To see why the last inequality holds, there are two cases to consider: (i) If $\max(I^{\prime})=\max(I)$ then the inequality follows immediately since $t_{i^{\prime}}\geqslant t_{m}$. (ii) If $\max(I^{\prime})<\max(I)$ then $i^{\prime}=\max(I)$ and $\max(I^{\prime})\geqslant\min(I^{\prime})=m$, so $t_{\max(I^{\prime})}+t_{i^{\prime}}-t_{m}\geqslant t_{i^{\prime}}=t_{\max(I)}$.
Now, let us show that $G-X$ does not contain $T_{i}$ as a minor, for some $i\in I$. By the inductive hypothesis, there exists some $i\in I^{\prime}$ such that $G^{\prime}-X^{\prime}$ does not contain a $T_{i}$ minor. Since $T_{i}$ is connected, any inclusion-wise minimal subgraph of $G$ containing a $T_{i}$ minor is connected. Since $B_{\ell}\subseteq X$, any such subgraph $M$ is contained in $G^{\prime}$ or is contained in $G_{\ell}-B_{\ell}$ or contains a vertex of $B_{\ell}$. In the first case, $M$ contains vertices of $X^{\prime}\subseteq X$, by the choice of $i$. The second case is ruled out by property ($\star$2). In the third case, $M$ contains a vertex of $B_{\ell}\subseteq X$. Thus, every subgraph of $G$ containing a $T_{i}$ minor contains a vertex of $X$, and hence $G-X$ contains no $T_{i}$ minor. This concludes the proof. ∎
We may now turn to the proof of Corollary 3. We will use the following lemma, which is a special case of a more general result of Robertson and Seymour [Statement (8.7) in [RS86]].
Lemma 5.
For every graph $G$, for every path decomposition $(B_{1},B_{2},\dots,B_{q})$ of $G$, for every family $\mathcal{F}$ of connected subgraphs of $G$, for every positive integer $d$, either:
  1. (1)
    there are $d$ pairwise vertex-disjoint subgraphs in $\mathcal{F}$, or
  2. (2)
    there is a set $X$ that is the union of at most $d-1$ bags of $(B_{1},B_{2},\dots,B_{q})$ such that $V(F)\cap X\neq\varnothing$ for every $F\in\mathcal{F}$.
Proof of Corollary 3.
It is known (and an easy exercise to show) that, for every positive integer $p$, the complete ternary tree $T_{p}$ of height $p$ has pathwidth $p$. First, apply Theorem 1 on $G$ with the tree $T_{p}$. If $G$ contains $k$ vertex-disjoint subgraphs each containing a $T_{p}$ minor, we are done. So we may assume that the theorem produces a set $X_{1}$ of at most $|V(T_{p})|(k-1)\leqslant 3^{p+1}(k-1)$ vertices such that $G-X_{1}$ has no $T_{p}$ minor.
By Lemma 4, $G-X_{1}$ has a path decomposition $(B_{1},B_{2},\dots,B_{q})$ of width strictly less than $3^{p+1}$. It is easily checked that every inclusion-wise minimal subgraph of $G-X_{1}$ with pathwidth at least $p$ is connected. Apply Lemma 5 on $G-X_{1}$ with the path decomposition $(B_{1},B_{2},\dots,B_{q})$, with $d=k$, and with the family $\mathcal{F}$ of connected subgraphs of $G-X_{1}$ with pathwidth at least $p$. If $\mathcal{F}$ contains $k$ pairwise vertex-disjoint members, we are done. So we may assume that the lemma produces a set $X_{2}$ of at most $3^{p+1}(k-1)$ vertices such that $X_{2}$ hits every member of $\mathcal{F}$. It follows that $G-X_{1}-X_{2}$ has pathwidth strictly less than $p$. Let $X\coloneqq X_{1}\cup X_{2}$. Since $|X|\leqslant 3^{p+1}(k-1)+3^{p+1}(k-1)\leqslant 2\cdot 3^{p+1}k$, the set $X$ has the desired properties. ∎
Acknowledgments
This work was done during a visit of Gwenaël Joret and Piotr Micek to the University of Ottawa and Carleton University. The research stay was partially funded by a grant from the University of Ottawa.